OpenStudy (anonymous):

Find the product of the complex number and its conjugate. 3 + 3i

OpenStudy (anonymous):

Given: 3 + 3i Conjugate: 3 - 3i So: (3 + 3i)(3 - 3i)

OpenStudy (anonymous):

If one complex no. is 3+3i then other has to be 3-3i So their product, that is (3+3i)(3-3i)= (3^2)-((3i)^2) The square of 3 will be 9 And the square of 3i will be -9 ( as i is square root of -1. So square of i will give us -1 which when multiplied by the square of 3 gives us -9) So the final answer will be 9-(-9)=18 :)

OpenStudy (anonymous):

thank u guys!!

OpenStudy (anonymous):

You're Welcome. :)